# Rotatable and Movable Antenna-Enabled Near-Field Integrated Sensing and Communication

### v2

@@ II-C Signal Model
where $\mathbf{W}=[\mathbf{w}_{1},\dots,\mathbf{w}_{K}]\in\mathbb{C}^{N_{\text{t}}\times K}$ denotes the digital beamformer for conveying the information symbol $\mathbf{s}(t)=\left[s_{1}(t),\dots,s_{K}(t)\right]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ to the UTs and $\mathbf{x}_{0}(t)\in\mathbb{C}^{N_{\text{t}}\times 1}$ denotes the dedicate sensing signal to achieve the full DoFs for target sensing [24]. The multiple beam transmission is also exploited by the dedicated sensing signal, hence its covariance matrix $\mathbf{R}_{0}=\mathbb{E}\left[\mathbf{x}_{0}(t)\mathbf{x}^{\mathsf{H}}_{0}(t)\right]$ is of a general rank. Moreover, the information symbols are modeled as independent Gaussian random variables with zero mean and unit power while the dedicated sensing signal is generated by pseudo-random coding, so that $\mathbb{E}\left[\mathbf{s}(t)\mathbf{s}^{\mathsf{H}}(t)\right]=\mathbf{I}_{K}$ and $\mathbb{E}\left[\mathbf{s}(t)\mathbf{x}^{\mathsf{H}}_{0}(t)\right]=\mathbf{0}_{K\times N_{\text{t}}}$. Thus the covariance matrix of the transmit signal $\mathbf{x}(t)$ is given by
@@
It is worth noting that $\mathbf{R}_{x}$ can be approximately calculated by
@@
where $\mathbf{G}=\eta\mathbf{g}_{0}\mathbf{h}_{0}^{\mathsf{T}}$ with $\eta$ denoting the complex channel gain of the sensing target, ${\mathbf{g}_{k}}=[{h}^{\text{r}}_{1,k};\dots,{h}^{\text{r}}_{N_{\text{t}},k}]^{\mathsf{T}}$, and $\mathbf{n}_{0}(t)\sim\mathcal{CN}(\mathbf{0}_{N_{\text{r}}},\sigma_{0}^{2}\mathbf{I}_{N_{\text{r}}})$ denotes the AWGN.
@@ II-D Sensing Performance Metric: CRB
CRB is a common metric for assessing sensing performance, which delineates a lower bound on the variance of unbiased estimators [25]. In this paper, we resort to the CRB as the performance metric for the estimation of the target’s positional information, i.e., the distance $d_{0}$, elevation angle $\theta_{0}$, and azimuth angle $\phi_{0}$. Denote $\bm{\zeta}=[d_{0},\theta_{0},\phi_{0},\eta^{\text{r}},\eta^{\text{i}}]$ as the vector of unknown parameters, where $\eta^{\text{r}}=\Re\{\eta\}$ and $\eta^{\text{r}}=\Im\{\eta\}$. Then the Fisher information matrix (FIM) for estimating $\bm{\zeta}$ can be partitioned as
@@
where $\mathbf{J}_{11}\in\mathbb{C}^{3\times 3}$ denotes the partition merely about $d_{0},\theta_{0}$, and $\phi_{0}$; $\mathbf{J}_{22}\in\mathbb{C}^{2\times 2}$ denotes the partition merely about $\eta$; and $\mathbf{J}_{12}\in\mathbb{C}^{3\times 2}$ denotes the partition about the mutual information between the position information and $\eta$. The detailed expressions of $\mathbf{J}_{11},\mathbf{J}_{22}$, and $\mathbf{J}_{12}$ are derived in Appendix VI. According to Woodbury matrix identity [26], the CRB matrix for estimating $d_{0},\theta_{0}$, and $\phi_{0}$ is given by
@@ II-E Communication Performance Metric: Communication Rate
In this paper, we assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism. From (20), the received SINR at the $k$-th UT can be expressed as
@@ III-A Problem Formulation
In this section, we consider a sensing-centric design by optimizing the sensing performance and holding a certain communication performance, which can be modeled as a CRB-minimization problem with communication rate constraints. Due to the monotonic increment of logarithmic function, the communication rate constraint for each user can be considered as a SINR constraint. Consequently, the corresponding SINR-constrained CRB-minimization problem is formulated as
@@
where $\mathbf{Q}^{\text{p}}=[\mathbf{q}_{1},\dots,\mathbf{q}_{N_{\text{p}}}]$; $P_{\max}\geq 0$ is the total transmit power budget; and $\Gamma_{\min}\geq 0$ represents the minimum SINR for the $k$-th user. Since $\mathcal{P}_{1}$ is non-convex and the variables are highly coupled, we adopt an AO-based framework to optimize the covariance matrix of the ISAC signal $\mathbf{R}_{x}$ as well as the RMAs’ positions and rotations in the following.
@@ III-B Optimization of Transmit Beamforming
Note that $\mathcal{P}_{3}$ is a convex semidefinite programming (SDP) problem, the global optimum of which can be obtained by the existing convex optimization solvers such as CVX and MOSEK. Although we omit the rank-one constraint of $\bm{\Omega}_{k}$, the rank-one global optimal $\mathbf{w}_{k}$ and corresponding $\mathbf{R}_{x}$ of $\mathcal{P}_{2}$ can always be constructed from an arbitrary global optimal solution of $\mathcal{P}_{3}$ by [24]:
@@
$\mathbf{R}_{x}^{\circ}=\tilde{\mathbf{R}}_{x},$ (33)
where $\tilde{\bm{\Omega}}_{k}$ and $\tilde{\mathbf{R}}_{x}$ are the optimal solution of $\mathcal{P}_{3}$ and optimal $\mathbf{R}_{0}$ can be obtained by $\mathbf{R}_{0}^{\circ}=\mathbf{R}_{x}^{\circ}-\sum_{k=1}^{K}\mathbf{w}_{k}^{\circ}(\mathbf{w}_{k}^{\circ})^{\mathsf{H}}$.
@@ III-C Optimization of RMAs’ Positions and Rotations
In this subsection, we optimize the RMAs’ positions and rotations, i.e., $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$, with $\mathbf{R}_{x}$ fixed. Since the objective function is highly non-convex and the solution space is large in general, it is difficult to obtain a local optimal solution. To efficiently address this challenge, PSO is introduced as an effective approach [27, 28]. In the PSO-based approach, we first initialize $M$ particles as
@@
Next, we update each particle according to its individual experience and the swarm experience, which can be described by the current local best particle $\bm{\xi}_{m,{\text{l}best}}$ and the current global best particle $\bm{\xi}_{\text{g}best}$. For the $\tau$-th iteration, the $m$-th particle is updated as:
@@
where $a_{1}$ and $a_{2}$ are the individual and global learning factors, which represent the step size of each particle moving toward the best position; $b_{1},b_{2}\sim\mathcal{U}(0,1)$ are two random parameters for enhancing the randomness of the search in order to escape from local optimum; $\omega$ is the parameter for maintaining the inertia of each particle’s movement. Specifically, we update $\omega$ by
@@
where $\omega_{\max}$ and $\omega_{\min}$ are the maximum and minimum values of $\omega$, respectively, with $\tau_{\max}$ denoting the maximum iteration number of PSO. Due to the fact that the particle may move out of the feasible region, we employ a projection function to process the updated particles in (36) as
@@
where ${\text{CRB}}(\bm{\xi}_{m}^{(\tau)})$ denotes the CRB calculated by setting RMAs’ positions and rotations according to $\bm{\xi}_{m}^{(\tau)}$; $\mu_{1},\mu_{2}$, and $\mu_{3}$ are the positive penalty parameters which are large enough; $\mathcal{S}(\bm{\xi})$ is a set composed of the position pairs of MAs that violate the minimum-distance constraint. Therefore, $\mathcal{S}(\bm{\xi})$ is defined as
@@
Similarly, $\mathcal{J}(\bm{\xi}_{m}^{(\tau)})$ is a set corresponding to the rotation constraints (7) and (8), which is defined as
$\mathcal{J}(\bm{\xi}_{m}^{(\tau)})$ =$=\left\{{n_{\text{r}}}|\mathbf{u}^{\text{t}}(\bm{\xi}^{(\tau)})\left(\mathbf{r}_{n_{\text{r}}}(\bm{\xi}^{(\tau)})-\mathbf{C}^{\text{t}}\right)>0,n_{\text{r}}\in[1,N_{\text{r}}]\right\}$
@@
where ${\mathbf{u}^{\text{p}}(\bm{\xi}^{(\tau)})}$, ${\mathbf{t}_{n_{\text{t}}}(\bm{\xi}^{(\tau)})}$, and $\mathbf{r}_{n_{\text{r}}}(\bm{\xi}^{(\tau)})$ represent the normal vector of the TP or RP, the $n_{\text{t}}$-th transmit MA’s position, and the $n_{\text{r}}$-th receive MA’s position calculated based on $\bm{\xi}^{(\tau)}$. Hence $|\mathcal{S}(\bm{\xi})|$ and $|\mathcal{J}(\bm{\xi}_{m}^{(\tau)})|$ are two adaptive penalty terms [29]. And $\mathcal{H}(\bm{\xi}_{m}^{(\tau)})$ is a brick wall penalty term corresponding to the SINR constraint (27d), which is defined as
@@ IV-A Problem Formulation
In this section, we consider a communication-centric design by optimizing the communication performance and holding a certain sensing performance, which can be modeled as a sum-rate-maximization problem with CRB constraint. Specifically, the corresponding CRB-constraint sum-rate-maximization problem is formulated as
@@ IV-B Problem Reformulation
where $\bm{\rho}=[\rho_{1},\dots,\rho_{K}]^{\mathsf{T}}$ and $f_{1}(\mathbf{W},\mathbf{R}_{0},\mathbf{H},\bm{\rho})=\sum_{k}^{K}\log_{2}(1+\rho_{k})-\rho_{k}+\frac{(1+\rho_{k})|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}|^{2}}{\sum_{i=1}|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{i}|^{2}+\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{0}\mathbf{h}_{k}^{\mathsf{*}}+\sigma_{k}^{2}}$. It is obvious that ${\mathcal{P}}_{\text{5}}$ is a convex problem w.r.t. ${\rho_{k}}$. Hence, by checking the first-order optimality condition, the optimal ${\rho_{k}}$ can be obtained as follows:
@@
where $\bm{\nu}=[\nu_{1},\dots,\nu_{K}]^{\mathsf{T}}$ denotes the auxiliary variables and $f_{2}(\mathbf{W},\mathbf{R}_{0},\mathbf{H},\bm{\rho},\bm{\nu})=\sum_{k=1}^{K}2(1+\rho_{k})\nu_{k}\sqrt{|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}|^{2}}-(1+\rho_{k})\nu_{k}^{2}(\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{x}\mathbf{h}_{k}^{\mathsf{*}}+\sigma_{k}^{2})$. Similarly, the optimal $\nu_{k}$ can be obtained by checking the first-order optimality condition as
@@
Thus, the original problem is reformulated as:
${\mathcal{P}}_{\text{7}}:$ $~{}\max_{\mathbf{W},\mathbf{R}_{0},\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}},\bm{\varpi}^{\text{r}}}~{}\sum_{k=1}^{K}2(1+\rho_{k})\nu_{k}\sqrt{\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}\mathbf{w}_{k}^{\mathsf{H}}\mathbf{h}_{k}^{*}}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad-(1+\rho_{k})\nu_{k}^{2}(\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{x}\mathbf{h}_{k}^{\mathsf{*}}+\sigma_{k}^{2})$ (49)

### v3

@@ II-C Signal Model
where $\mathbf{W}=[\mathbf{w}_{1},\dots,\mathbf{w}_{K}]\in\mathbb{C}^{N_{\text{t}}\times K}$ denotes the digital beamformer for conveying the information symbol $\mathbf{s}(t)=\left[s_{1}(t),\dots,s_{K}(t)\right]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ to the UTs and $\mathbf{x}_{0}(t)\in\mathbb{C}^{N_{\text{t}}\times 1}$ denotes the dedicate sensing signal to achieve the full DoFs for target sensing [25]. The multiple beam transmission is also exploited by the dedicated sensing signal, whose covariance matrix $\mathbf{R}_{0}=\mathbb{E}\left[\mathbf{x}_{0}(t)\mathbf{x}^{\mathsf{H}}_{0}(t)\right]$ is of a general rank. Moreover, the information symbols are modeled as independent Gaussian random variables with zero mean and unit power while the dedicated sensing signal is generated by pseudo-random coding, so that $\mathbb{E}\left[\mathbf{s}(t)\mathbf{s}^{\mathsf{H}}(t)\right]=\mathbf{I}_{K}$ and $\mathbb{E}\left[\mathbf{s}(t)\mathbf{x}^{\mathsf{H}}_{0}(t)\right]=\mathbf{0}_{K\times N_{\text{t}}}$. Thus the covariance matrix of the transmit signal $\mathbf{x}(t)$ is given by
@@
As per [26], $\mathbf{R}_{x}$ can be approximately calculated by
@@
where $\mathbf{G}=\eta\mathbf{g}_{0}\mathbf{h}_{0}^{\mathsf{T}}$ with $\eta$ denoting the complex channel gain of the sensing target, ${\mathbf{g}_{0}}=[{h}^{\text{r}}_{1,0};\dots,{h}^{\text{r}}_{N_{\text{t}},0}]^{\mathsf{T}}$, and $\mathbf{n}_{0}(t)\sim\mathcal{CN}(\mathbf{0}_{N_{\text{r}}},\sigma_{0}^{2}\mathbf{I}_{N_{\text{r}}})$ denotes the AWGN.
@@ II-D Sensing Performance Metric: CRB
In the near-field region, the spherical wave propagation enables joint estimation of distance and angles of the target, which can be observed in (20). In this paper, we aim to estimate the target’s positional information, i.e., the distance $d_{0}$, elevation angle $\theta_{0}$, and azimuth angle $\phi_{0}$ and resort to the CRB as the performance metric since CRB delineates a lower bound on the variance of unbiased estimators [27]. Denote $\bm{\zeta}=[d_{0},\theta_{0},\phi_{0},\eta^{\text{r}},\eta^{\text{i}}]$ as the vector of unknown parameters, where $\eta^{\text{r}}=\Re\{\eta\}$ and $\eta^{\text{i}}=\Im\{\eta\}$. Then the Fisher information matrix (FIM) for estimating $\bm{\zeta}$ can be partitioned as
@@
where $\mathbf{J}_{11}\in\mathbb{C}^{3\times 3}$ denotes the partition merely about $d_{0},\theta_{0}$, and $\phi_{0}$; $\mathbf{J}_{22}\in\mathbb{C}^{2\times 2}$ denotes the partition merely about $\eta$; and $\mathbf{J}_{12}\in\mathbb{C}^{3\times 2}$ denotes the partition about the mutual information between the position information and $\eta$. The detailed expressions of $\mathbf{J}_{11},\mathbf{J}_{22}$, and $\mathbf{J}_{12}$ are derived in Appendix -A. According to Woodbury matrix identity [28], the CRB matrix for estimating $d_{0},\theta_{0}$, and $\phi_{0}$, which is the $3\times 3$ submatrix in the upper-left corner of $\mathbf{J}_{\bm{\zeta}}^{-1}$, is given by
@@ II-E Communication Performance Metric: Communication Rate
In this paper, we assume that the perfect channel state information (CSI) of communication channels is perfectly known at the BS via a proper channel estimation mechanism [29]. From (24), the received SINR at the $k$-th UT can be expressed as follows:
@@ III-A Problem Formulation
In this section, we consider a sensing-centric design by optimizing the sensing performance while guaranteeing a certain communication performance, which can be modeled as a CRB-minimization problem with communication rate constraints. Due to the monotonic increment of logarithmic function, the communication rate constraint for each user can be considered as a SINR constraint. Consequently, the corresponding SINR-constrained CRB-minimization problem is formulated as
@@
where $\mathbf{Q}^{\text{p}}=[\mathbf{q}_{1},\dots,\mathbf{q}_{N_{\text{p}}}]$; $P_{\max}\geq 0$ is the total transmit power budget; and $\Gamma_{\min}\geq 0$ represents the minimum SINR for the $k$-th user. Since $\mathcal{P}_{1}$ is non-convex and the variables are highly coupled, we adopt an AO-based framework to optimize ISAC signal’s covariance matrix $\mathbf{R}_{x}$ as well as the RMAs’ positions and rotations in the following.
@@ III-B Optimization of Transmit Beamforming
Note that $\mathcal{P}_{3}$ is a convex semidefinite programming (SDP) problem, the global optimum of which can be obtained by the existing convex optimization solvers such as CVX and MOSEK. Although we omit the rank-one constraint of $\bm{\Omega}_{k}$, the rank-one global optimal $\mathbf{w}_{k}$ and corresponding $\mathbf{R}_{x}$ of $\mathcal{P}_{2}$ can always be constructed from an arbitrary global optimal solution of $\mathcal{P}_{3}$.
Proposition 1.
Given an arbitrary global optimum $\tilde{\mathbf{R}}_{x}$, $\tilde{\bm{\Omega}}_{1},\dots,\tilde{\bm{\Omega}}_{K}$ of $\mathcal{P}_{3}$, the following solution is a global optimum of $\mathcal{P}_{2}$:
@@
$\mathbf{R}_{x}^{\circ}=\tilde{\mathbf{R}}_{x}.$ (37)
Proof:
Please refer to Appendix -B. ∎
Then optimal $\mathbf{R}_{0}$ can be obtained by $\mathbf{R}_{0}^{\circ}=\mathbf{R}_{x}^{\circ}-\sum_{k=1}^{K}\mathbf{w}_{k}^{\circ}(\mathbf{w}_{k}^{\circ})^{\mathsf{H}}$.
@@ III-C Optimization of RMAs’ Positions and Rotations
In this subsection, we optimize the RMAs’ positions and rotations, i.e., $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$, with $\mathbf{R}_{x}$ fixed. Since the objective function is highly non-convex and the solution space is large in general, it is difficult to obtain a local optimal solution. To efficiently address this challenge, PSO is introduced as an effective approach [30, 31]. In the PSO-based approach, we first initialize $M$ particles as follows:
@@
Next, we update each particle according to its individual experience and the swarm experience, which can be described by the current local best particle $\bm{\xi}_{m,{\text{l}best}}$ and the current global best particle $\bm{\xi}_{\text{g}best}$. For the $\tau$-th iteration, the $m$-th particle is updated as follows:
@@
where $a_{1}$ and $a_{2}$ are the individual and global learning factors, which represent the step size of each particle moving toward the best position; $b_{1},b_{2}\sim\mathcal{U}(0,1)$ are two random parameters for enhancing the randomness of the search in order to escape from local optimum; $\omega$ is the parameter for maintaining the inertia of each particle’s movement. Specifically, we update $\omega$ as follows:
@@
where $\omega_{\max}$ and $\omega_{\min}$ are the maximum and minimum values of $\omega$, respectively, with $\tau_{\max}$ denoting the maximum iteration number of PSO. Due to the fact that the particle may move out of the feasible region, we employ a projection function to process the updated particles in (40) as follows:
@@
where ${\text{CRB}}(\bm{\xi}_{m}^{(\tau)})$ denotes the CRB calculated by setting RMAs’ positions and rotations according to $\bm{\xi}_{m}^{(\tau)}$; $\mu_{1},\mu_{2}$, and $\mu_{3}$ are the positive penalty parameters which are large enough; $\mathcal{S}(\bm{\xi})$ is a set composed of the position pairs of MAs that violate the minimum-distance constraint. Specifically, $\mathcal{S}(\bm{\xi})$ is defined as follows:
@@
Similarly, $\mathcal{J}(\bm{\xi}_{m}^{(\tau)})$ is a set corresponding to the rotation constraints (11) and (12), which is defined as follows:
$\mathcal{J}(\bm{\xi}_{m}^{(\tau)})=\left\{{n_{\text{r}}}|\mathbf{u}^{\text{t}}(\bm{\xi}^{(\tau)})\left(\mathbf{r}_{n_{\text{r}}}(\bm{\xi}^{(\tau)})-\mathbf{C}^{\text{t}}\right)>0,n_{\text{r}}\in[1,N_{\text{r}}]\right\}$
@@
where ${\mathbf{u}^{\text{p}}(\bm{\xi}^{(\tau)})}$, ${\mathbf{t}_{n_{\text{t}}}(\bm{\xi}^{(\tau)})}$, and $\mathbf{r}_{n_{\text{r}}}(\bm{\xi}^{(\tau)})$ represent the normal vector of the TP or RP, the $n_{\text{t}}$-th transmit MA’s position, and the $n_{\text{r}}$-th receive MA’s position calculated based on $\bm{\xi}^{(\tau)}$. Hence $|\mathcal{S}(\bm{\xi})|$ and $|\mathcal{J}(\bm{\xi}_{m}^{(\tau)})|$ are two adaptive penalty terms [32]. And $\mathcal{H}(\bm{\xi}_{m}^{(\tau)})$ is a brick wall penalty term corresponding to the SINR constraint (31d), which is defined as follows:
@@ IV-A Problem Formulation
In this section, we consider a communication-centric design by optimizing the communication performance while ensuring a certain sensing performance, which can be modeled as a sum-rate-maximization problem with CRB constraint. Specifically, the corresponding CRB-constraint sum-rate-maximization problem is formulated as
@@ IV-B Problem Reformulation
where $\bm{\rho}=[\rho_{1},\dots,\rho_{K}]^{\mathsf{T}}$ and $f_{1}(\mathbf{W},\mathbf{R}_{0},\mathbf{H},\bm{\rho})=\sum_{k}^{K}\left[\log_{2}(1+\rho_{k})-\rho_{k}+\frac{(1+\rho_{k})|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}|^{2}}{\sum_{i=1}|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{i}|^{2}+\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{0}\mathbf{h}_{k}^{\mathsf{*}}+\sigma_{k}^{2}}\right]$. It is obvious that ${\mathcal{P}}_{\text{5}}$ is a convex problem w.r.t. ${\rho_{k}}$. Hence, by checking the first-order optimality condition, the optimal ${\rho_{k}}$ can be obtained as follows:
@@
where $\bm{\nu}=[\nu_{1},\dots,\nu_{K}]^{\mathsf{T}}$ denotes the auxiliary variables and $f_{2}(\mathbf{W},\mathbf{R}_{0},\mathbf{H},\bm{\rho},\bm{\nu})=\sum_{k=1}^{K}[2(1+\rho_{k})\nu_{k}\sqrt{|\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}|^{2}}-(1+\rho_{k})\nu_{k}^{2}(\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{x}\mathbf{h}_{k}^{\mathsf{*}}+\sigma_{k}^{2})]$. Similarly, the optimal $\nu_{k}$ can be obtained by checking the first-order optimality condition as
@@
Thus, the original problem is equivalently reformulated as:
${\mathcal{P}}_{\text{7}}:$ $~{}\max_{\mathbf{W},\mathbf{R}_{0},\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}},\bm{\varpi}^{\text{r}}}~{}\sum_{k=1}^{K}\left[2(1+\rho_{k})\nu_{k}\sqrt{\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}\mathbf{w}_{k}^{\mathsf{H}}\mathbf{h}_{k}^{*}}\right.$
[…]